OpenStudy (anonymous):

x + 2y - z = -1 x - 4y - z = -4 2x - 2y + 3z = 15 solve using matrices

OpenStudy (anonymous):

@mathmath333 i think you helped yesterday but my computer went down. if you could help again thatd be great :)

OpenStudy (phi):

can you write down the matrix ? i.e. order the variables alphabetical (easier to remember what is what if you do that, then write the coefficients in matrix form what do you get ?

OpenStudy (anonymous):

1 2 -1 x -1 1 -4 -1 X y = -4 2 -2 3 z 15

OpenStudy (phi):

ok, but I would leave out the multiply sign (it looks like an "x". One way to solve is find the inverse of the matrix (tedious for a 3 x 3) another is to use "row reduction" (not too bad) What techniques have you been studying ?

OpenStudy (anonymous):

matrices but i dont understand it

OpenStudy (phi):

yes, but there are a few different ways. *any* idea how they want you to solve this ? If not, we can use row reduction.

OpenStudy (anonymous):

no specific way

OpenStudy (phi):

ok, then let's use row reduction. to start, can you answer this. IF you had the matrix 1 2 -1 x 9 -1 1 -4 y = 8 -0 0 1 z 2 do you know what z is ?

OpenStudy (anonymous):

no i dont

OpenStudy (phi):

to answer, change the last row "back" into the original equation (remember the numbers are the coefficients for x y and z )

OpenStudy (anonymous):

how would you change it back?

OpenStudy (phi):

how did you write the matrix? you "ignored" the variable names, and just wrote the numbers example, the first equation x + 2y - z = -1 became the first row: 1 2 -1 -1 (I left out the unknown)

OpenStudy (phi):

to go back, put in the x, y and z in the last row of my made-up matrix: 0 0 1 2

OpenStudy (anonymous):

x + y + z = 2 ?

OpenStudy (phi):

that would be if had the row 1 1 1 2 but you have 0 0 1 2

OpenStudy (anonymous):

0x + 0y + 1z = 2

OpenStudy (phi):

yes, and we can simplify that. zero times anything is zero.

OpenStudy (anonymous):

ok

OpenStudy (phi):

in other words, you have no x's and no y's

OpenStudy (phi):

and get z= 2

OpenStudy (anonymous):

oh ok. so how would i do that with the equations up top

OpenStudy (phi):

that is the idea with "reduction". It's a way to change the matrix so we have a bunch of zeros except in one place. (Like the last row of my example)

OpenStudy (anonymous):

oh. i get your example but not mine

OpenStudy (phi):

yes, but now we can tackle your problem 1 2 -1 -1 1 -4 -1 -4 2 -2 3 15 the idea is we want to change the "leading number" (the first number) in each row (except the first row) into zero. to do this we multiply the top row by a number so that its first number (the 1) when added to the second row, gives us zero. so off to the side, we say: the second row has a 1 that we want to change to zero. We do this by adding -1. I need to multiply the top row by -1 to change the first number into -1.

OpenStudy (phi):

in other words, we multiply the top row by -1 (that means multiply each number in the top row by -1. we get -1 * (1 2 -1 -1) --> -1 -2 +1 + 1

OpenStudy (phi):

now add -1 -2 +1 + 1 to the second row (number by number) can you write down what you get ?

OpenStudy (anonymous):

0 -2 -1 -3

OpenStudy (phi):

Here is the original matrix 1 2 -1 -1 1 -4 -1 -4 2 -2 3 15 Off to the side we multiply the first row by -1 and get -1 -2 +1 +1 add that to the *second row*: -1 -2 +1 +1 1 -4 -1 -4 -------------- <--- sum goes here

OpenStudy (phi):

can you try again?

OpenStudy (anonymous):

0 -6 0 -3

OpenStudy (phi):

yes. that becomes our "new" second row. the problem is now 1 2 -1 -1 0 -6 0 -3 2 -2 3 15 Now let's change the first 2 in the last row to a zero. what do we add to 2 to make it 0?

OpenStudy (anonymous):

add -2

OpenStudy (phi):

yes. what to we multiply the first row by so it's first number becomes -2 ?

OpenStudy (anonymous):

1

OpenStudy (phi):

1 times the first row will give you 1 2 -1 -1 you want the -2 ...rest of row... in other words, we need to multiply the 1 by something to get -2 if it's not clear, you can say 1*x = -2 what is x ?

OpenStudy (anonymous):

-2

OpenStudy (phi):

though I should not use "x". I mean what unknown number N should we use? Yes -2 can you write down what you get when you multiply the first row by -2?

OpenStudy (anonymous):

the first row being 1 2 -1 ?

OpenStudy (phi):

yes, but with the other -1 (on the other side of the = sign) 1 2 -1 -1 I know this is a bit mysterious...

OpenStudy (anonymous):

-2 -4 2 2

OpenStudy (phi):

yes, now add that to the third row in our matrix: 1 2 -1 -1 0 -6 0 -3 2 -2 3 15

OpenStudy (anonymous):

0 -6 5 17

OpenStudy (phi):

with row reduction we would next tackle the next column in and zero out the entries below the the second row. But in this case notice the second row already has two zeros

OpenStudy (anonymous):

i see

OpenStudy (phi):

can you write down the equation to "goes with" the second row ?

OpenStudy (anonymous):

-6y = -3

OpenStudy (phi):

can you solve for y ?

OpenStudy (anonymous):

y=.5

OpenStudy (phi):

yes. next we tackle the bottom row which has only y and z in it. write down the equation that you get from the third row.

OpenStudy (anonymous):

-6y + 5z = 17

OpenStudy (phi):

and we know y= 1/ 2 put in 1 /2 for y, and solve for z. can you do that ?

OpenStudy (anonymous):

i dont understand

OpenStudy (phi):

you have -6y + 5z = 17 we already figured out y is 1/2 so replace y with 1/2 -6 * 1/2 + 5z = 17 simplify and solve for z. do you know how ?

OpenStudy (phi):

as a first step, I would multiply -6 * 0.5

OpenStudy (anonymous):

i got 3.3

OpenStudy (anonymous):

@phi

OpenStudy (phi):

can you show your work?

OpenStudy (phi):

-6 * 1/2 + 5z = 17 becomes -3 + 5z = 17 next, add +3 to both sides.

OpenStudy (anonymous):

1/2 + 5z = 17 add - 1/2 to both sides 5z = 17 - 1/2 5z = 16.5 divide both sides by 5 and get 3.3

OpenStudy (phi):

Oh. you did 1y + 5z = 17 and replaced y with 1/2 to get 1/2 + 5z = 17 the work looks good. but we are doing the last row, which is 0 -6 5 17 and that becomes -6*y + 5 z = 17 with y=1/2 it becomes -6* 1/2 + 5z = 17 now solve for z

OpenStudy (anonymous):

can you help me im not sure how to do it

OpenStudy (phi):

when you see two numbers being multiplied, it's a good idea to simplify to get one number -6* 1/2 + 5z = 17

OpenStudy (anonymous):

-3 + 5z = 17 5z = 17 + 3 5z = 20 z= 4

OpenStudy (phi):

most excellent. we now know y = 1/2 and z = 4 now we use the top row. care to tackle it ?

OpenStudy (anonymous):

-6y + 5z = 17?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

we already used that row we need to solve for "x" and we use the first row in the matrix

OpenStudy (anonymous):

ohh so -2x + -4y + 2z = 2

OpenStudy (phi):

yes, except I noticed I pasted in the wrong matrix. It should be 1 2 -1 -1 0 -6 0 -3 0 -6 5 17 we used the second and third rows to find y and then z

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so does that change the equation

OpenStudy (phi):

we have to use the correct matrix. The one I pasted with -2 -4 2 2 was just *wrong* we want to use 1 2 -1 -1 0 -6 0 -3 0 -6 5 17

OpenStudy (anonymous):

ok

OpenStudy (phi):

so we use 1 2 -1 -1 can you find "x" ?

OpenStudy (anonymous):

x+2y -z =-1

OpenStudy (phi):

yes. and now replace y with 1/2 and z with 4

OpenStudy (anonymous):

-5.5

OpenStudy (phi):

go slower. one step at a time. first, *just* replace y and z with their numbers in x+2y -z =-1 what do you get ?

OpenStudy (anonymous):

x + 1/2 - 4=-1

OpenStudy (phi):

ok but you should not leave out the coefficients. when you start with x+2y -z =-1 you replace the y with 1/2 but *you keep the 2*

OpenStudy (anonymous):

oh okay

OpenStudy (phi):

in other words, the equation x + 2y - z = -1 says x plus *TWO* y's take away z is -1 we need to multiply the y by 2

OpenStudy (anonymous):

qnd get 4y?

OpenStudy (phi):

I meant when we start with x + 2y - z = -1 and replace y with 1/2 we write x + 2*1/2 - z = -1 and replace z with 4: x + 2* 1/2 - 4 = -1 ok?

OpenStudy (anonymous):

ok

OpenStudy (phi):

can you finish ? x + 2* 1/2 - 4 = -1

OpenStudy (anonymous):

X=2

OpenStudy (phi):

yes. we have solved the problem if you multiply the original matrix times the vector 2 0.5 4 you will get the right hand side vector -1 -4 15

OpenStudy (phi):

also, each of the 3 original equations will be true if we use the values x=2, y=1/2 and z=4

OpenStudy (phi):

If you have time, Khan goes into detail how to do this type of problem https://www.khanacademy.org/math/precalculus/precalc-matrices/reduced_row_echelon/v/matrices-reduced-row-echelon-form-1